​(Comprehensive EOQ calculations​) Knutson Products Inc. is involved in the production of airplane parts and has the following​ inventory, carrying, and storage​ costs:

1. Orders must be placed in round lots of 100 units.

2. Annual unit usage is 300,000. ​(Assume a​ 50-week year in your​ calculations.)

3. The carrying cost is 30 percent of the purchase price.

4. The purchase price is ​$30 per unit.

5. The ordering cost is $500 per order.

6. The desired safety stock is 6,000 units.​ (This does not include​ delivery-time stock.)

7. The delivery time is 3 week.

Given the foregoing​ information:

a. Determine the optimal EOQ level.

b. How many orders will be placed​ annually?

c. What is the inventory order​ point? (That​ is, at what level of inventory should a new order be​ placed?)

d. What is the average inventory​ level?

e. What would happen to the EOQ if annual unit sales doubled​ (all other unit costs and safety stocks remaining​ constant)? What is the elasticity of EOQ with respect to​ sales? (That​ is, what is the percentage change in EOQ divided by the percentage change in​ sales?)

f. If carrying costs​ double, what will happen to the EOQ ​level? (Assume the original sales level of 300,000 ​units.) What is the elasticity of EOQ with respect to carrying​ costs?

g. If the ordering costs​ double, what will happen to the level of EOQ​? ​(Again, assume original levels of sales and carrying​ costs.) What is the elasticity of EOQ with respect to ordering​ costs?

h. If the selling price​ doubles, what will happen to EOQ​?

What is the elasticity of EOQ with respect to selling​ price?
Answer #1

Solution:

a)Calculation of EOQ

EOQ=SQRT[2*Annual demand*Cost per order/Carrying cost]

=SQRT[2*300,000*$500/($30*30%)]

=5773.5 or 5774

b)No. of orders that can be placed

=Annual demand/EOQ

=300,000/5774

=52 orders.

c)Inventory reorder level

No.of days in a year=50*7=350 days

Reorder Level=(Average weekly usage*Average lead time)+safety stock

=[(300,000/50)*3]+6000

=24000 units

d)Average inventory level=EOQ/2

=5774/2

=2887 units